# DCAD-2000: A Multilingual Dataset across 2000+ Languages with Data Cleaning as Anomaly Detection

### v2

@@ 3.2.2 Anomaly Detection
$\tilde{x}_{j}=\frac{x_{j}-\mu_{j}}{\sigma_{j}},\quad j=1,\ldots,8,$ (2)
@@
$\tilde{\mathbf{x}}=\frac{\mathbf{x}-\boldsymbol{\mu}}{\boldsymbol{\sigma}},$ (3)
@@
Take Isolation Forest Liu et al. (2008) as an example [^10] We also evaluate some other algorithms, please refer to Section 5 for more details., we compute an anomaly score $\phi(\tilde{\mathbf{x}})$ for each document. The Isolation Forest algorithm assigns anomaly scores based on the average path length required to isolate a data point in a decision tree. Specifically, for a document represented by $\tilde{\mathbf{x}}$, the anomaly score is defined as:
$\phi(\tilde{\mathbf{x}})=2^{-\frac{h(\tilde{\mathbf{x}})}{c(n)}},$ (4)
where $h(\tilde{\mathbf{x}})$ is the average path length for $\tilde{\mathbf{x}}$ across all trees in the Isolation Forest, and $c(n)$ is the average path length of a point in a binary tree with $n$ samples, given by:
$c(n)=2H(n-1)-\frac{2(n-1)}{n},$ (5)
@@
$f(\tilde{\mathbf{x}})=\begin{cases}1,&\text{if }\phi(\tilde{\mathbf{x}})<\tau,\\ -1,&\text{if }\phi(\tilde{\mathbf{x}})\geq\tau,\end{cases}$ (6)
@@
Once the anomaly scores $\phi(\tilde{\mathbf{x}})$ are computed for all samples in the standardized dataset $\tilde{\mathcal{X}}=\{\tilde{\mathbf{x}}_{1},\ldots,\tilde{\mathbf{x}}_{N}\}$, we partition the dataset into two subsets:
$\mathcal{X}_{\text{keep}}$ =$=\{\tilde{\mathbf{x}}\in\tilde{\mathcal{X}}:f(\tilde{\mathbf{x}})=1\},$ (7)
$\mathcal{X}_{\text{remove}}$ =$=\{\tilde{\mathbf{x}}\in\tilde{\mathcal{X}}:f(\tilde{\mathbf{x}})=-1\}.$ (8)
@@ 3.2.3 Visualization
To qualitatively assess the separation achieved by our anomaly detection framework, we generate scatter plots (Figure 1) of the $8$ feature values, with data points color-coded according to their anomaly labels. These visualizations facilitate the interpretation of decision boundaries and highlight the features that contribute most significantly to the detection process. We observe well-defined clusters separating anomalous and non-anomalous data points, with anomalies exhibiting distinct patterns compared to the majority of the data. Features such as the language identification score ($s_{\text{lid}}(t)$) and perplexity score ($s_{\text{ppl}}(t)$) are expected to be particularly discriminative in identifying anomalies, as they capture linguistic irregularities and unexpected text patterns. For example, low $lid$ or unusually high $ppl$ scores often indicate problematic text, such as spam, low-quality content, or noise. The framework effectively identifies and removes such low-quality text samples, which can be easily visualized by the separation of these points in the scatter plots.
@@ 3.3 Computational Resources
The construction of the DCAD-2000 dataset leveraged clould servers to process and clean the multilingual data efficiently. Each server instance is equipped with 32 CPU cores, 128GB of memory, and 100GB of disk storage, which is utilized for intermediate data handling and memory-intensive operations such as anomaly detection. The workload is managed using container orchestration tools, with up to 100 parallel tasks running per job to ensure scalability.
@@ 4 Dataset Analysis
Document Distribution Across Data Sources. The DCAD-2000 dataset is derived from four primary sources: MaLA, Fineweb, Fineweb-2, and Newly Extracted Common Crawl data (New CC), as described in Section 3.1. Figure 2(a) presents the distribution of documents across these sources, with Fineweb-2 and New CC collectively contributing 47.5% and 39.3% of the total dataset, respectively. These two sources play a significant role in ensuring the dataset’s emphasis on both language diversity (Fineweb-2) and corpus freshness (New CC). MaLA, though contributing 11.1% of the total dataset, brings in valuable content from non-Common Crawl sources, further enriching the diversity of the dataset, especially for low-resource languages and specialized domains.
Geographical Coverage of Languages. The geographical distribution of languages in DCAD-2000, based on Glottolog’s classification [^12] Geographic data source: https://glottolog.org, is shown in Figure 2(b). The dataset spans languages from all major world regions, with the largest proportions originating from Africa (28.6%), Papunesia (26.3%) and Eurasia (23.8%). This coverage ensures robust support for multilingual applications across varied regional contexts, including densely populated areas like Eurasia and sparsely populated regions such as Papunesia and Australia. While Eurasia is more heavily represented, this diversity of linguistic coverage helps ensure that the dataset remains useful for training LLMs in diverse regional environments.
@@
Language Resource Classification. Following the classification approach proposed by Goyal et al. (2022), we categorize languages in DCAD-2000 into four groups based on corpus size: high-resource, medium-resource, low-resource, and extremely low-resource. Table 1 shows the distribution across these categories. The dataset includes 155 high- and medium-resource languages, while low-resource languages make up a significant portion, which reflects DCAD-2000’s commitment to supporting underrepresented languages. Notably, DCAD-2000 surpasses other corpora in its balance between high-resource and low-resource languages, which can have a significant impact on multilingual model training. The distribution of languages across categories ensures that the dataset is well-suited for developing models that perform effectively across diverse language resources.
Impact of Data Cleaning. We summarize the document count, token count, and disk size of the high/medium/low resource languages in DCAD-2000 before and after the data cleaning process. Complete details are provided in Appendix C. The cleaning process results in the removal of a substantial amount of noisy data, even from datasets like MaLA, Fineweb, and Fineweb-2, which had already been subject some cleaning. This aligns with the findings from Dou et al. (2024); Zhang et al. (2024). For example, in the MaLA dataset, 8.05 million documents are removed for the hbs_Latn language, which suggests the necessity of rigorous data cleaning to enhance dataset quality. Overall, the cleaning process removed approximately 7.69% of the documents across all languages, significantly improving the quality of the dataset by reducing noise and increasing relevance for model training.
@@ 5 Evaluation
To validate the effectiveness of the proposed data cleaning method and the quality of the DCAD-2000 dataset, we conduct a series of experiments on FineTask benchmark [^13] https://huggingface.co/spaces/HuggingFaceFW/blogpost-fine-tasks. FineTask includes tasks in nine languages: Chinese, French, Arabic, Russian, Thai, Hindi, Turkish, Swahili, and Telugu, covering diverse NLP tasks such as reading comprehension, common-sense reasoning, natural language understanding, and text generation. We continue pretrain the LLaMA-3.2-1B model Dubey et al. (2024) in these nine languages and analyze the impact of different data cleaning methods, anomaly detection algorithms, and existing datasets. We use normalized accuracy as the evaluation metric. Detailed experimental settings and results are provided in Appendix D.
Baseline Heuristic Threshold Our
Arabic 0.07 0.17 0.21
Turkish 0.07 0.22 0.27
Swahili 0.08 0.13 0.29
Russian 0.10 0.19 0.24
Telugu 0.02 0.06 0.06
Thai 0.14 0.13 0.18
Chinese 0.12 0.18 0.32
French 0.11 0.23 0.35
Hindi 0.07 0.15 0.21
Table 2: The performance comparison of models trained using various data cleaning methods.
Impact of Different Data Cleaning Strategies. We compare the performance of models trained using various data cleaning strategies to evaluate the effectiveness of our proposed anomaly detection-based cleaning framework. As shown in Table 2, the baseline method (raw, uncleaned data) consistently underperforms all cleaning strategies, demonstrating the critical importance of data quality in multilingual training. This poor performance can be attributed to the presence of noisy, irrelevant, or inconsistent data, which hampers model generalization. Traditional threshold-based filtering [^14] We use the implementation from https://github.com/bigscience-workshop/data-preparation applies fixed rules based on features such as language identification scores and document complexity to remove low-quality data. In contrast, our anomaly detection-based cleaning approach dynamically identifies and filters noisy or anomalous data, resulting in significantly better model performance. Models trained on cleaned data using our framework achieved a normalized accuracy improvement of approximately 5–20% over the baseline and outperformed the threshold-based method by 3–10%. In terms of scalability, while the threshold-based filtering method is computationally efficient due to its simple, rule-based nature, our anomaly detection-based approach requires more computation. However, it is highly effective at capturing nuanced data issues that fixed rules might miss, making it a better choice for high-quality multilingual training in the long term.
Baseline Iso_Forest OC_SVM LOF K-Means
Arabic 0.07 0.21 0.18 0.21 0.14
Turkish 0.07 0.27 0.29 0.17 0.15
Swahili 0.08 0.29 0.25 0.19 0.19
Russian 0.10 0.24 0.19 0.18 0.15
Telugu 0.02 0.06 0.05 0.04 0.04
Thai 0.14 0.21 0.18 0.18 0.15
Chinese 0.12 0.32 0.28 0.25 0.21
French 0.11 0.35 0.37 0.30 0.23
Hindi 0.07 0.21 0.17 0.16 0.14
@@
Comparison of Anomaly Detection Algorithms. We evaluate several classical anomaly detection algorithms to determine the optimal method for constructing DCAD-2000. The evaluated detection approaches include Isolation Forest (ISO_Forest; Liu et al., 2008), One-Class SVM (OC_SVM; Manevitz and Yousef, 2001), Local Outlier Factor (LOF; Breunig et al., 2000), and K-Means Hartigan et al. (1979), using implementations from scikit-learn. [^15] https://scikit-learn.org Table 3 illustrates the performance of these algorithms in cleaning the dataset. While all anomaly detection-based methods outperform the baseline, the performance of OC_SVM, LOF, and K-Means is inconsistent. These methods often require extensive fine-tuning of parameters like the number of neighbors (for LOF) or the kernel choice (for OC_SVM), which significantly affects results and increases computational overhead. In contrast, ISO_Forest demonstrates more stable and robust performance across experiments, largely due to its ability to efficiently handle noisy and high-dimensional multilingual data. Unlike other methods, ISO_Forest does not require intensive hyperparameter tuning to achieve reliable results, making it ideal for large-scale multilingual datasets. However, there are trade-offs with ISO_Forest: it can be computationally expensive compared to simpler algorithms like K-means, especially when working with very high-dimensional data (Our feature vectors has 8 dimentions as mentioned in Section 3.2.2). Nevertheless, its robustness and scalability make it the most suitable choice for cleaning multilingual data in DCAD-2000.
Figure 3: Compare DCAD-2000 with other multilingual corpora.
@@
Comparison with Other Multilingual Datasets. To validate the quality of DCAD-2000, we compare it against existing multilingual corpora on the FineTask benchmark. These corpora include datasets constructed from New CC, MaLA, and Fineweb-2 as described in Section 3.1. As shown in Figure 3, models trained on DCAD-2000 consistently outperform those trained on other datasets, achieving higher normalized accuracy. The improvements can be attributed to the enhanced data quality, diversity, and reduced noise resulting from our comprehensive cleaning pipeline. Specifically, DCAD-2000 provides greater linguistic diversity and a more balanced representation of low-resource languages, leading to improved performance on tasks involving underrepresented languages like Swahili and Telugu. This demonstrates that DCAD-2000 not only offers a stronger foundation for multilingual NLP research but also makes it directly suitable for training large-scale multilingual models.

### v3

@@ 3.2.2 Anomaly Detection
$\tilde{x}_{j}=\frac{x_{j}-\mu_{j}}{\sigma_{j}},\quad j=1,\ldots,8.$ (2)
@@
$\tilde{\mathbf{x}}=\frac{\mathbf{x}-\boldsymbol{\mu}}{\boldsymbol{\sigma}}.$ (3)
@@
[…]
$\phi(\tilde{\mathbf{x}})=2^{-\frac{h(\tilde{\mathbf{x}})}{c(n)}}.$ (4)
[…]
$c(n)=2H(n-1)-\frac{2(n-1)}{n}.$ (5)
@@
[…]
@@
[…]
$\mathcal{X}_{\text{keep}}$ $=\{\tilde{\mathbf{x}}\in\tilde{\mathcal{X}}:f(\tilde{\mathbf{x}})=1\},$ (7)
$\mathcal{X}_{\text{remove}}$ $=\{\tilde{\mathbf{x}}\in\tilde{\mathcal{X}}:f(\tilde{\mathbf{x}})=-1\}.$ (8)
@@ 3.2.3 Visualization
To qualitatively evaluate the separation achieved by our data cleaning framework, we present scatter plots of the eight feature dimensions in Figure 1, with data points color-coded by their anomaly labels. These visualizations facilitate the interpretation of decision boundaries and highlight the features that contribute most significantly to the detection process. We observe well-defined clusters separating anomalous and non-anomalous data points, with anomalies exhibiting distinct patterns compared to the majority of the data. Features such as the language identification score ($s_{\text{lid}}(t)$) and perplexity score ($s_{\text{ppl}}(t)$) are expected to be particularly discriminative in identifying anomalies, as they capture linguistic irregularities and unexpected text patterns. For example, low lidlid or unusually high pplppl scores often indicate problematic text, such as spam, low-quality content, or noise. The framework effectively identifies and removes such low-quality text samples, which can be easily visualized by the separation of these points in the scatter plots.
@@ 3.3 Computational Resources
The construction of the DCAD-2000 dataset leveraged Ksyun servers [^10] https://k8s.ksyun.com to process and clean the multilingual data efficiently. Each server instance is equipped with 32 CPU cores, 128GB of memory, and 100GB of disk storage, which is utilized for intermediate data handling and memory-intensive operations such as anomaly detection. The workload is managed using container orchestration tools, Kubernetes [^11] https://kubernetes.io, with up to 100 parallel tasks running per job to ensure scalability.
@@ 4 Dataset Analysis
Document Distribution Across Data Sources. The DCAD-2000 dataset is derived from four primary sources: MaLA, Fineweb, Fineweb-2, and Newly Extracted Common Crawl data (New CC), as described in Section 3.1. Figure 2(a) presents the distribution of documents across these sources, with Fineweb-2 and New CC collectively contributing 47.5% and 39.3% of the total dataset, respectively. These two sources play a significant role in ensuring the dataset’s emphasis on both language diversity (Fineweb-2) and corpus freshness (New CC). MaLA, though contributing 11.1% of the total dataset, brings in valuable content from non-Common Crawl sources, further enriching the diversity of the dataset, especially for low-resource languages.
Geographical Coverage of Languages. Figure 2(b) shows the geographical distribution of languages in DCAD-2000, based on the number of unique languages available in each region, as classified by Glottolog. [^12] Geographic data source: https://glottolog.org The dataset spans languages from all major world regions, with the largest proportions originating from Africa (28.6%), Papunesia (26.3%) and Eurasia (23.8%). This coverage ensures robust support for multilingual applications across varied regional contexts, including densely populated areas like Eurasia and sparsely populated regions such as Papunesia and Australia. While Eurasia is more heavily represented, this diversity of linguistic coverage helps ensure that the dataset remains useful for training LLMs in diverse regional environments.
@@
Language Resource Classification. Following the classification approach proposed by Flores [17], we categorize languages in DCAD-2000 into four groups based on corpus size: high-resource, medium-resource, low-resource, and extremely low-resource. Table 1 shows the distribution across these categories. The dataset includes 155 high- and medium-resource languages, while low-resource languages make up a significant portion, which reflects DCAD-2000’s commitment to supporting underrepresented languages. Notably, DCAD-2000 surpasses other corpora in its balance between high-resource and low-resource languages, which can have a significant impact on multilingual model training. The distribution of languages across categories ensures that the dataset is well-suited for developing models that perform effectively across diverse language resources.
Impact of Data Cleaning. We summarize the document count, token count, and disk size of the high/medium/low resource languages in DCAD-2000 before and after the data cleaning process. Complete details are provided in Appendix C. The cleaning process results in the removal of a substantial amount of noisy data, even from datasets like MaLA, Fineweb, and Fineweb-2, which had already been subject some cleaning. This aligns with the findings from [18, 19]. For example, in the MaLA dataset, 8.05 million documents are removed for the hbs_Latn language, which suggests the necessity of rigorous data cleaning to enhance dataset quality. Overall, the cleaning process removed approximately 7.69% of the documents across all languages, significantly improving the quality of the dataset by reducing noise and increasing relevance for model training (Section 5).
@@ 5 Evaluation
Following Fineweb-2 [14], we conduct a series of experiments on the FineTask benchmark [^13] https://huggingface.co/spaces/HuggingFaceFW/blogpost-fine-tasks to evaluate the effectiveness of our proposed data cleaning pipeline and assess the quality of the DCAD-2000 dataset. FineTask comprises tasks in nine languages—Chinese, French, Arabic, Russian, Thai, Hindi, Turkish, Swahili, and Telugu—and covers a diverse set of NLP tasks, including reading comprehension, commonsense reasoning, natural language understanding, and text generation. To investigate the impact of different data cleaning strategies and anomaly detection algorithms, we continue pretraining on three typical LLMs: LLaMA-3.2-1B[47], Qwen-2.5-7B[48], and Aya-expanse-32B[49]. Additionally, we analyze the performance across different resource categories using the SIB-200[24], Glot500-c [8], and FLORES-200 [25] benchmarks. We report normalized accuracy for FineTask, raw accuracy for SIB-200, negative log-likelihood (NLL) for Glot500-c, and BLEU scores for FLORES-200. Full experimental settings and results are provided in Appendix D.
Figure 3: The performance comparison of models trained using various data cleaning methods.
Impact of Different Data Cleaning Strategies. We evaluate the effectiveness of our proposed anomaly detection-based data cleaning framework by comparing model performance across various cleaning strategies. As illustrated in Figure 3, the baseline model trained on raw, unfiltered data consistently underperforms relative to all cleaning methods. This performance gap is primarily due to noisy, irrelevant, or inconsistent data that hinders model generalization. Traditional threshold-based filtering [^14] We use the implementation from https://github.com/bigscience-workshop/data-preparation, which removes low-quality samples using fixed rules based on features, yields modest improvements. In contrast, our anomaly detection-based approach dynamically identifies and filters anomalous or noisy data, resulting in significantly enhanced model performance. Models trained using our method achieve normalized accuracy improvements of approximately 5–20% over the baseline, and outperform the threshold-based approach by 3–10%. While the threshold-based method is more computationally efficient due to its simplicity, our framework—though more resource-intensive—effectively captures subtle data issues that static rules often overlook.
@@
LLaMA-3.2-1B Qwen-2.5-7B Aya-expanse-32B
Baseline Iso_Forest OC_SVM LOF K-Means Baseline Iso_Forest OC_SVM LOF K-Means Baseline Iso_Forest OC_SVM LOF K-Means
Arabic 0.07 0.21 0.18 0.21 0.14 0.63 0.71 0.68 0.65 0.68 0.69 0.75 0.70 0.71 0.69
Turkish 0.07 0.27 0.29 0.17 0.15 0.65 0.72 0.73 0.67 0.68 0.75 0.79 0.77 0.76 0.77
Swahili 0.08 0.29 0.25 0.19 0.19 0.25 0.34 0.27 0.35 0.27 0.35 0.44 0.36 0.37 0.41
Russian 0.10 0.24 0.19 0.18 0.15 0.74 0.79 0.75 0.75 0.76 0.77 0.82 0.79 0.80 0.79
Telugu 0.02 0.06 0.05 0.04 0.04 0.16 0.24 0.26 0.20 0.21 0.15 0.25 0.19 0.21 0.27
Thai 0.14 0.21 0.18 0.18 0.15 0.57 0.64 0.59 0.59 0.61 0.38 0.46 0.42 0.43 0.40
Chinese 0.12 0.32 0.28 0.25 0.21 0.75 0.82 0.77 0.76 0.78 0.69 0.75 0.71 0.71 0.73
French 0.11 0.35 0.37 0.30 0.23 0.74 0.80 0.76 0.76 0.75 0.74 0.79 0.76 0.76 0.76
Hindi 0.07 0.21 0.17 0.16 0.14 0.49 0.57 0.52 0.53 0.52 0.69 0.74 0.72 0.73 0.72
@@
Comparison of Anomaly Detection Algorithms. We compare several classical anomaly detection algorithms to identify the most effective approach for constructing DCAD-2000. The evaluated methods include Isolation Forest (ISO_Forest;46), One-Class SVM (OC_SVM; 50), Local Outlier Factor (LOF;51), and K-Means [52], using implementations from scikit-learn [^15] https://scikit-learn.org. We provide the comparison of different algorithms in Appendix D.3. Table 2 reports the performance of these algorithms in cleaning the dataset. While all anomaly detection methods outperform the unfiltered baseline, the performance of OC_SVM, LOF, and K-Means is notably inconsistent. These algorithms often require extensive parameter tuning—such as selecting the number of neighbors for LOF or the kernel type for OC_SVM, which introduces sensitivity to hyperparameters and increases computational overhead. In contrast, ISO_Forest demonstrates more stable and robust performance across experiments, attributed to its efficiency in handling noisy, high-dimensional multilingual data. Unlike other methods, ISO_Forest delivers reliable results without intensive hyperparameter tuning, making it particularly suitable for large-scale multilingual datasets. However, ISO_Forest can be more computationally demanding than simpler methods like K-Means, especially in high-dimensional settings (our feature vectors have eight dimensions, as described in Section 3.2.2). Despite this trade-off, its robustness and scalability establish ISO_Forest as the most appropriate choice for data cleaning in DCAD-2000.
(a) French – LLaMA-3.2-1B
(b) Chinese – Qwen-2.5-7B
(c) Turkish – Aya-expanse-32B
Figure 4: Comparison of DCAD-2000 with existing multilingual corpora for three languages—French, Chinese, and Turkish—evaluated using different multilingual LLMs.
@@
Comparison with Other Multilingual Datasets. To validate the quality of DCAD-2000, we compare it against existing multilingual corpora on the FineTask benchmark. These corpora include datasets constructed from New CC, MaLA, and Fineweb-2 as described in Section 3.1. As shown in Figure 4, models trained on DCAD-2000 consistently outperform those trained on other datasets, achieving higher normalized accuracy. The improvements can be attributed to the enhanced data quality, diversity, and reduced noise resulting from our comprehensive cleaning pipeline. Specifically, DCAD-2000 provides greater linguistic diversity and a more balanced representation of low-resource languages, leading to improved performance on tasks involving underrepresented languages like Swahili and Telugu.
Table 3: Performance across different language categories. We use accuracy ($\uparrow$) in SIB-200, negative log-likelihood ($\downarrow$) in Glot500-c and BLEU ($\uparrow$) in FLORES-200. Improvements are highlighted accordingly.
LLaMA-3.2-1B Qwen-2.5-7B Aya-expanse-32B
Fineweb-2 New CC DCAD-200 Fineweb-2 New CC DCAD-200 Fineweb-2 New CC DCAD-200
SIB-200 ($\uparrow$) H 8.24 8.86 10.37 ↑2.13 33.41 34.53 38.26 ↑4.85 41.72 42.41 47.93 ↑6.21
M 7.31 7.92 9.15 ↑1.84 28.72 29.86 32.65 ↑3.93 32.25 33.39 38.16 ↑5.91
L 6.06 6.45 7.83 ↑1.77 23.58 24.22 27.12 ↑3.54 26.87 27.57 33.24 ↑6.37
VL 3.68 4.27 5.24 ↑1.56 13.25 15.43 21.57 ↑8.32 17.23 19.5 26.38 ↑9.15
Glot500-c test ($\downarrow$) H 426.37 403.58 373.14 ↓53.23 347.21 334.18 303.38 ↓43.83 273.85 257.24 225.28 ↓48.57
M 446.28 436.94 423.75 ↓22.53 385.72 389.24 369.15 ↓16.57 326.92 321.16 302.53 ↓24.39
L 503.38 493.27 473.96 ↓29.42 426.33 419.25 404.28 ↓22.05 372.62 367.26 341.34 ↓31.28
VL 584.55 569.34 532.86 ↓51.69 479.04 463.36 433.48 ↓45.56 396.33 392.33 385.86 ↓10.47
FLORES-200 ($\uparrow$)
(Eng–X)
H 3.14 3.82 5.26 ↑2.12 15.24 16.07 18.47 ↑3.23 23.45 24.33 26.33 ↑2.88
M 2.75 2.94 3.89 ↑1.14 12.83 13.46 15.49 ↑2.66 19.36 20.21 21.62 ↑2.26
L 2.27 2.41 3.14 ↑0.87 8.94 9.28 10.25 ↑1.31 16.61 17.24 18.36 ↑1.75
VL 1.85 2.05 2.35 ↑0.50 6.33 7.25 9.05 ↑2.72 12.51 13.16 14.77 ↑2.26
FLORES-200 ($\uparrow$)
(X–Eng)
H 3.94 3.98 4.26 ↑0.32 16.31 16.92 18.84 ↑2.53 23.86 24.13 26.94 ↑3.08
M 3.52 3.66 3.80 ↑0.28 13.65 14.05 16.27 ↑2.62 20.45 20.36 22.53 ↑2.17
L 3.05 3.12 3.24 ↑0.19 9.47 10.22 11.48 ↑2.01 17.67 17.82 18.93 ↑1.26
VL 2.73 2.83 3.14 ↑0.41 7.28 7.81 9.65 ↑2.37 13.25 13.56 15.88 ↑2.63
Analysis by the Categories of Language Resources. Table 3 presents model performance across languages categorized by resource levels (High, Medium, Low, and Very Low). Across all benchmarks and model sizes, DCAD-2000 consistently outperforms Fineweb-2 and New CC. While gains are modest in high-resource languages, improvements are substantial for low- and very low-resource languages—up to +9.15 accuracy on SIB-200 and -53.23 NLL on Glot500-c—highlighting the effectiveness of our cleaning pipeline in enhancing data quality where it is most needed. The BLEU results on FLORES-200 further validate these trends, with notable improvements in both English-to-X and X-to-English translation tasks. These consistent gains across tasks and languages demonstrate that DCAD-2000 enables more balanced multilingual performance and is well-suited for training inclusive, high-quality language models.